OpenStudy (anonymous):

on the opening night of a play at a local theatre, 959 tickets were sold for a total of $8176 adult tickets cost $10 each. children's tickets cost $8 each, and senior citizen tickets cost $6 each. If the combined number of children and adult tickets exceeded twice the number of senior citizen tickets by 218, then how many tickets of each type were sold?

OpenStudy (anonymous):

i have one answer correct which is 247 but my first and second answer is wrong i had 499 for x and 213 for y. what are the correct numbers?

OpenStudy (anonymous):

here we have 3 variables and 3 equations. this gives us a degree of freedom of zero. we can therfore solve

OpenStudy (anonymous):

the equations are: x+y+z=959 and 10x+8y+6z=8176 and x+y=2z+218. Can you see how i derived these from the info given? (x is adult y is children and z is senior)

OpenStudy (anonymous):

Not sure how youre instructor wants you to do it but that it how i would do it.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

oh i had them switched around

OpenStudy (anonymous):

So can you solve that system of 3 equations?

OpenStudy (anonymous):

yes i have the answer for equation 3